A man carries a load of 50 kg to a height of 15m in 5 minutes. Calculate A) the work done, B)the power extended
1
Expert's answer
2019-03-15T12:45:28-0400

a) By the definition of the work, we have:


"W = Fh = mgh = 50 kg \cdot 15 m \cdot 9.8 \dfrac{m}{s^2} = 7350 J."

b) We can find the power from the formula:


"P = \dfrac{W}{t} = \dfrac{7350 J}{5 min \cdot \dfrac{60 s}{1 min}} = 24.5 W."

Answer:

a) "W = 7350 J."

b) "P = 24.5 W."